OpenStudy (kkutie7):

Find the volume V of the described solid S. The base of S is a circular disk with radius 5r. Parallel cross-sections perpendicular to the base are squares.

OpenStudy (kkutie7):

A=pi(radius)^2 ?

OpenStudy (kkutie7):

this is calculus so i doubt its that easy

OpenStudy (rational):

a rough sketch might help draw a circle of given radius and try to visualize the cross sections perpendicular to this circle

OpenStudy (aum):



OpenStudy (rational):

the cross sections will be in another dimension that shoots out of this screen

OpenStudy (kkutie7):

I follow so far

OpenStudy (dumbcow):

@aum, isnt \[AB = 2 \sqrt{25r^2 -y^2}\]

OpenStudy (rational):

\[\large V = \int \limits_{-5}^5 \pi (2x)^2 dy\] \(\large x^2+y^2 = 5^2\)

OpenStudy (dumbcow):

@rational , there is no "pi" because its only looking at area of the square

OpenStudy (rational):

Ahh yeah will fix it :) \[\large V = \int \limits_{-5}^5 (2x)^2 dy = \int \limits_{-5}^5 \left( 2\sqrt{25-y^2}\right)^2 dy \]

OpenStudy (rational):

does that look okay, i deliberately removed r as i thought radus is 5 and not 5r

OpenStudy (kkutie7):

so your are using the radius as a sort of x values here from -5 to 5 and the function its the horizontal line of the cross-section?

OpenStudy (rational):

the volume element here is a square of very thin radus the side of the square is 2x and you can only see top view in below picture :